OpenStudy (anonymous):

1/10 divided by 6/7 reduced to lowest terms using a mixed # or a Whole #

OpenStudy (anonymous):

1/10 divided by 6/7 equals 1/10 * 7/6 = 7/60 which is in lowest terms

OpenStudy (anonymous):

when dividing by a fraction just multiply by the reciprocal

OpenStudy (anonymous):

eg 5/2=2.5 and 5*.5=2.5

OpenStudy (anonymous):

huh the answer was 3/35 i dont know how they reduced it 7 does not go into 3

OpenStudy (anonymous):

that would be the answer to 1/10 div 7/6

OpenStudy (anonymous):

1/10 * 6/7 = 6/70 = 3/35